The question says: Determine whether each ordered pair is a solution of y = -2x + 5.
1.) (3, -5)
2.) (2.5, 0)
Please explain how you got the answers. Thanks

Answer by algebrapro18(249)

All you have to do is plug in the x values for x and the y values for y and see if you get the same number.

y = -2x+5 (3,-5)
-5 = -2(3)+5
-5 = -6+5
-5 = -1 nope this is not a solution because -5 does not = -1.

y = -2x+5 (2.5,0)
0 = -2(2.5)+5
0 = -5 + 5
0 = 0 yes this is a solution, because 0 does equal 0.